OpenStudy (superhelp101):

Explain, in complete sentences, the relationship between radicals and complex numbers. Give an example illustrating this relationship.

OpenStudy (kohai):

In a radical equation, at least one variable of the equation is under a radical sign, or (equivalently) has a fractional exponent. Example: sqrt(x+1) = 25 Or (x+1)^(-1/2) = 25 which is the same equation, just expressed differently An equation with just a radical in it could have the radical (root) of a number or a constant involved in the terms of the equation. y^3 +sqrt(A) = sqrt(7) is an example of such an equation, since the variable y does not have a fractional power (equivalent to having a radical sign over it) Radicals is the name mathematics uses to define terms under a square root sign. Complex Numbers take the form a + bi, where a and b are Real Numbers while i represents the square root of -1 Sq.rt.(x+4) is a radical [3 + 4(i)] is a Complex Number Radicals become Complex Numbers whenever the term under the sq.rt. sign is negative. Thus, sq.rt.(-4) =sq.rt.(4)(sq.rt.(-1), =+/-2i

OpenStudy (superhelp101):

ohh okai. Thx!

OpenStudy (kohai):

You're welcome